4) Consider an agent with initial wealth W to be allocated between a safe (risk-free) asset and a risky asset. Let xs denote4.2) Solve for the agents optimal investment portfolio (xs and xr) assuming negative investments are not permissible.
Answer #1

The agent needs to solve the following maximization problem to obtain the optimal investment portfolio:

Maximize π. u((1+r​​​​​​r)x​​​​​​r+ x​​​​​​s )+ (1-π).u((1+r​​​​​​r)x​​​​​​r+ x​​​s )

= π.ln(x​​​​​​s ) + (1-π).ln(2xr + x​​​​​​s )

Now W= x​​​​​​s + x​​​​​​r

x​​​​​​r= W - x​​​​​​s . Now, substituting this in the agent's problem and taking first order derivative w.r.t x​​​​​​s , we get:

π/x​​​​​​s - (1-π)/(2W - x​​​​​​s ) = 0

We get x​​​​​​s = 2πW and x​​​​​​r = W(1-2π)